OpenStudy (symone19):

A shelf has 20 bags of potatoes. Two of the bags have potatoes starting to go bad. What is the expected probability of the complement of the event "bad bags of potatoes"? 0.1 0.2 0.8 0.9

OpenStudy (lfghjyt):

0.1

OpenStudy (brooke..help00):

.2? That is how many fell out of the bag.

OpenStudy (lfghjyt):

Wait nvm

OpenStudy (brooke..help00):

That are bad not how many fell out.

OpenStudy (lfghjyt):

There are 2 bad bags out of 20, so the probability of bad bags is 2/10 = 1/10 = 0.1 But we want the complement, which is the probability of bags which are not bad, and that's 1 - 0.1 = d] 0.9 as discussed above.

OpenStudy (lovelynerds):

0.1